OpenStudy (anonymous):

A rope parallel to the ground pulls a block across ice. The boulder has a mass of 500kg and there are three beefy guys wearing crampons exerting a constant net force of 2000N on the rope. What is the acceleration of the block? How long does it take the boulder to slide 5 meters to the edge of the pond?

OpenStudy (experimentx):

theoritically, F = m*a or, a = 2000/500 = 4 m/s^2 and if initial velocity is 0 solving this equation gives time 5 = 1/2*4*t^2 or t = sqrt(t) .... however physics don't work this way, it's impossible to apply constant net force of 2000N on a moving body, after it reaches the peak velocity, the net force would be zero ... (net force -friction = 0) so the body moves at constant velocity

OpenStudy (experimentx):

oops sorry .. i would like to take back "impossible" ... we generally don't do so